Each straight stretch of a 400-meter running track is 110 m long. Each curved part is 90 m long. Find the distance between the inside curbs of the two straight streches.
Answer by jorel1380(3719)

1/2 Circumference=90 m
1 Circumference=180
2πr=180
πr=90
r=90/π
d=2x(90/π)=(180/π)m..